# Rethinking Spectral Graph Neural Networks with Spatially Adaptive Filtering

### 17 Jan 2024 v1

@@ 4 Rethinking Spectral GNNs from the Spatial Perspective
4.1 Interplay of Spectral and Spatial Domains through the Lens of Graph Optimization
The graph signal denoising problem (Shuman et al., 2013) was initially leveraged in (Ma et al., 2021; Zhu et al., 2021) as a means to interpret GNNs with smoothness assumption, which yet does not always hold in certain real-world graph scenarios such as heterophily (Zhu et al., 2020a). Without loss of generality, in this paper, we consider a more generalized graph optimization problem  [^1] This problem was first introduced in (He et al., 2021) to derive theoretically grounded graph filters. However, in this study, we repurpose it as a bridge between spectral filtering and spatial aggregation.
$\mathop{\arg\min}_{\mathbf{Z}}\ \mathcal{L}=\alpha\|\mathbf{X}-\mathbf{Z}\|^{2}_{F}+(1-\alpha)\cdot\mathrm{tr}(\mathbf{Z}^{T}\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z})$ (3)
@@ 4.1 Interplay of Spectral and Spatial Domains through the Lens of Graph Optimization
$\mathbf{Z}^{*}=(\mathbf{I}+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\hat{\mathbf{L}}))^{-1}\mathbf{X}=\mathbf{U}(\mathbf{I}+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\mathbf{\Lambda}))^{-1}\mathbf{U}^{T}\mathbf{X}=\mathbf{U}g_{\psi}(\mathbf{\Lambda})\mathbf{U}^{T}\mathbf{X}.$ (4)
@@
$\mathbf{Z}^{(k)}=\alpha\mathbf{X}+(1-\alpha)\hat{\mathbf{A}}^{\text{new}}\mathbf{Z}^{(k-1)},\quad k=1,2,\cdots,K$ (5)
Theoretical Interaction — the Adapted New Graph. With the non-negative constraint, it is evident that both spectral filtering and spatial aggregation effectively address the generalized graph optimization problem in Eq. (3), despite their distinctive forms and operation domains. Upon closer examination, we discover a compelling relationship between the graph filter $g_{\psi}(\lambda)$ in Eq. (4) and the new graph $\hat{\mathbf{A}}^{\text{new}}$ in Eq. (5), given $g_{\psi}(\lambda)=(1+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\mathbf{\lambda}))^{-1}$, that is,
$\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\gamma_{\theta}(\hat{\mathbf{L}})=\mathbf{I}-\frac{\alpha}{1-\alpha}(g_{\psi}(\hat{\mathbf{L}})^{-1}-\mathbf{I})$ (6)
@@
It is worthy noting that this new graph fundamentally differs from the graph operation $g_{\psi}(\hat{\mathbf{L}})$ used in spectral filtering. Whereas the former elucidates the inherent spatial node relationships, the latter primarily processes node features without a clear physical interpretation. Moreover, it is crucial to understand that $g_{\psi}(\hat{\mathbf{L}})$ may not be fully-connected, especially with fixed-order polynomial approximation. This is because it only captures up to a $K$-hop neighborhood, i.e., $g_{\psi}(\hat{\mathbf{L}})=\sum_{k=0}^{K}\psi_{k}P_{k}(\hat{\mathbf{L}})=\sum_{k=0}^{K}\omega_{k}\hat{\mathbf{A}}^{k}$. In contrast, this uncovered new graph intrinsically enjoys a non-local property, as confirmed in the following section. Building upon this inherent non-locality, we further devise a framework to overcome the limitations of the truncated polynomials (see details in Section 5).
Figure 1: Distributions of adjacent nodes in the new graph based on their geodesic distance in the original graph. It is seen that nodes, distant in the original graph, can be connected in the new graph.
4.2 In-depth Analysis of the adapted new graph
@@
With the established non-negative constraint on graph filters, specifically $0<g_{\psi}(\lambda)\leq 1$, it becomes evident that the eigenvalues of $\mathbf{I}-g_{\psi}(\hat{\mathbf{L}})$ falls into the interval permitting Neumann series expansion, as shown in lemma 1 (Horn & Johnson, 2012). Building on this observation, we present a non-trivial property of the new graph in the following proposition (see proof in Appendix B.1).
Proposition 1.
[…]
This proposition implies that the new graph engenders immediate links between nodes that originally necessitate multiple hops for connection. To further underpin this theoretical claim, we analyze the general connection status on the new graph by BernNet (He et al., 2021), a spectral GNN adhering to the non-negative constraint. From Figure 1, it is apparent that nodes originally separated by multiple hops achieve direct connections in the new graph. See more empirical evidences in Appendix F.1.
[…]
4.2.2 Signed Edge Weights — Discerning Label Consistency.
Upon further scrutinizing the adapted new graph, we make a notable discovery that it readily accommodates both positive and negative edge weights. A more granular analysis in Figure 2 reveals that a considerable portion of positive edge weights are assigned to the same-class node pairs, enhancing graph homophily (see the formula in Appendix D). Conversely, edges parameterized with negative weights tend to bridge nodes with different labels. These findings demonstrate the newfound graph’s adeptness in discerning label consistency among nodes. To theoretically explain this phenomenon, we further present the proposition below (see proof in Appendix B.2):
Proposition 2.
[…]
@@
5 Spatially Adaptive Filtering Framework
Building on our discoveries, we re-evaluate the state-of-the-art spectral GNNs and put forth a paradigm-shifting framework, Spatially Adaptive Filtering (SAF), for joint exploitation of graph-structured data across both spectral and spatial domains. SAF leverages the adapted new graph by spectral filtering for an auxiliary non-local aggregation, addressing enduring challenges in GNNs related to long-range dependencies and graph heterophily. See the overall pipeline in Appendix A.
Non-negative Spectral Filtering. The proposed SAF requires explicit computation of the newfound graph, as outlined in Eq. (6). This further necessitates the graph filter $g_{\psi}:\left[0,2\right]\to\mathbb{R}$ to satisfy the non-negative constraint from Eq. (3): $0\leq g_{\psi}(\lambda)\leq 1$. However, not all extant graph filters fulfill this prerequisite. For instance, the filter use by GCN (Kipf & Welling, 2017), $g_{\psi}(\lambda)=1-\lambda$, takes negative values when $\lambda>1$. In this research, we approximate the graph filter using Bernstein polynomials (Farouki, 2012), which are known for their non-negative traits (Powers & Reznick, 2000) and are essential in a preeminent spectral GNN, BernNet (He et al., 2021). For $g_{\psi}(\lambda)\leq 1$ part, we rescale Bernstein polynomials with the following proposition.
Proposition 3.
@@ 5 Spatially Adaptive Filtering Framework
where $f_{\varphi}(\cdot)$, a two-layer MLP, maps $\mathbf{X}$ from $f$ to $C$ dimensions using 64 hidden units, and $\{\psi_{k}\}_{k=1}^{K}$ are non-negative learnable parameters. See Appendix G for alternative spectral filters in SAF.
@@
Non-local Spatial Aggregation. Once acquiring a suitable spectral filter $\hat{g}_{\psi}(\lambda)$, we promptly compute the adapted new graph using Eq. (6), represented as $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\tau(\mathbf{U}g_{\psi}(\mathbf{\Lambda})^{-1}\mathbf{U}^{T}-\mathbf{I})$ where $\tau=\frac{\alpha}{1-\alpha}$ serves as a scaling hyper-parameter and eigendecomposition is utilized for efficient matrix inversion.
Equipped with this newfound graph, we proceed to perform non-local aggregation:
$\mathbf{Z}^{(l)}=(1-\eta)\mathbf{Z}^{(0)}+\eta\hat{\mathbf{A}}^{\text{new}}\mathbf{Z}^{(l-1)},\quad l=1,2,\cdots,L,$
@@
where $\eta$ refers to the update rate and $\mathbf{Z}^{(0)}=f_{\varphi}(\mathbf{X})$. The iteratively aggregated results are denoted as $\mathbf{Z}_{a}$. Recognizing the potential noise from the non-local nature of $\hat{\mathbf{A}}^{\text{new}}$, we employ a sparsification technique, leveraging a positive threshold $\epsilon$, and retain only essential elements outside the $\left[-\epsilon,\epsilon\right]$ interval. For clarity, this refined model is referred to as SAF-$\epsilon$.
@@
Node-wise Prediction Amalgamation. To optimally use information from different graph domains, our method employs an attention mechanism, allowing nodes to determine the importance of each space. This mechanism produces pairwise weights for a nuanced amalgamation during model prediction. Specifically, the weight pair is computed as $\boldsymbol{\kappa}_{f}=\text{Sigmoid}(\mathcal{P}_{f}(\mathbf{Z}_{f})),\boldsymbol{\kappa}_{a}=\text{Sigmoid}(\mathcal{P}_{a}(\mathbf{Z}_{a}))$ where $\mathcal{P}_{f}(\cdot)$ and $\mathcal{P}_{a}(\cdot)$ are two different mapping functions that transforms elements from the domain $\mathbb{R}^{C}$ to $\mathbb{R}$, and $\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\in\mathbb{R}^{N}$ contain the weights for each node. For simplicity, we implement them using two one-layer MLPs. Given domain predictions $\mathbf{Y}_{f},\mathbf{Y}_{a}\in\mathbb{R}^{C}$, the final model prediction is attained as
$\mathbf{Y}=\mathrm{diag}(\boldsymbol{\kappa}_{f})\cdot\mathbf{Y}_{f}+\mathrm{diag}(\boldsymbol{\kappa}_{a})\cdot\mathbf{Y}_{a}\ \text{~{}where~{}}\ \left[\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\right]\leftarrow\frac{\left[\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\right]}{\max\{\|\{\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\}\|_{1},\delta\}}.$ (7)
@@
Here, $\delta$ is a small value to prevent non-zero division, with $\boldsymbol{\kappa}_{f}+\boldsymbol{\kappa}_{a}=\mathbf{1}$ maintained. Similar learning scheme can be founded in works (Sagi & Rokach, 2018; Wang et al., 2020; Zhu et al., 2020b).
Complexity. The proposed SAF augments current spectral GNNs with non-local aggregation and node-wise amalgamation. It requires a precomputation step of eigendecomposition which, despite involving the complexity of $\mathcal{O}(N^{3})$, can be practically expedited (Lanczos, 1950; Cai et al., 2021) and efficiently reused for both training and inference. Readers are directed to Appendix C.3 for further remarks on the contemporary significance of eigendecomposition. Non-local aggregation entails creating a new graph and propagating information over it. In SAF-$\epsilon$, these two steps are distinctly separated by a sparsification procedure, culminating in a complexity of $\mathcal{O}(N^{3}+N^{2}+nnz(\hat{\mathbf{A}}^{\text{new}})d)$ where $nnz$ denotes non-zero element count. Conversely, the standard SAF, viewing non-local aggregation holistically, can reduce complexity to $\mathcal{O}(2dN^{2}+dN)$ when $d\ll N$. For node-wise amalgamation, its parallelizable nature ensures computational efficiency. We also present empirical studies on both time and space overheads in Appendix F.5.
6 Experiments
Datasets We evaluate models over 13 real-world datasets from various domains. These include three well-known homophilic graphs: Cora, Citeseer, and Pubmed (Sen et al., 2008), five commonly used heterophilic graphs: Chameleon, Squirrel (Rozemberczki et al., 2021), Cornell, Texas (Pei et al., 2020), and Actor (Tang et al., 2009), as well as five recently introduced benchmarks: Twitch-DE (Lim et al., 2021) (social network), Minesweeper (synthetic graph), Tolokers (crowdsourcing platform worker network), Amazon-ratings (product co-purchasing), and Roman-empire (word dependency graph) (Platonov et al., 2023). We list their statistics in Appendix D.
Baselines. (1) MLP; (2) basic GNNs: GCN (Kipf & Welling, 2017) and APPNP (Gasteiger et al., 2019); (3) SOTA spatial GNNs: GCNII (Chen et al., 2020), FAGCN (Bo et al., 2021), GEN (Wang et al., 2021), PDE-GCN (Eliasof et al., 2021), NodeFormer (Wu et al., 2022) and GloGNN++ (Li et al., 2022); (4) SOTA spectral GNNs: ARMA (Bianchi et al., 2021), GPR-GNN (Chien et al., 2021), BernNet (He et al., 2021), ChebNetII (He et al., 2022), JacobiConv (Wang & Zhang, 2022), Specformer (Bo et al., 2023), LON-GNN (Tao et al., 2023) and OptBasisGNN (Guo & Wei, 2023a); (5) Unified GNNs: GNN-LF (Zhu et al., 2021), GNN-HF (Zhu et al., 2021), ADA-UGNN (Ma et al., 2021) and FE-GNN (Sun et al., 2023). Models like Geom-GCN (Pei et al., 2020) and Non-Local GNNs (Liu et al., 2021), surpassed by these SOTA methods, are excluded from our comparison.
Setup. To follow (He et al., 2021; 2022; Wang & Zhang, 2022), we fix $K=10$. For each dataset, we perform a grid search to tune the hyper-parameters of all models. With the best hyper-parameters, we train models with Adam optimizer (Kingma & Ba, 2014) in 1,000 epochs using early-stopping strategy and a patience of 200 epochs, and report the mean classification accuracies with a 95% confidence interval on 10 random data splits. More experimental details can be found in Appendix E.
@@ 6 Experiments
Table 1: Semi-supervised node classification accuracy (%) $\pm$ 95% confidence interval.
@@
Improv. 14.56% 9.68% 15.37% 13.99% 3.97% 1.62% 1.86% 0.68%
Semi-supervised Node Classification. In this task, we follow the experimental protocol established by (He et al., 2022) and compare our models with MLP, two basic GNNs, seven popular polynomial spectral GNNs, and four unified GNNs. For data splitting on homophilic graphs (Cora, Citeseer, and Pubmed), we apply the standard division (Yang et al., 2016) with 20 nodes per class for training, 500 nodes for validation, and 1,000 nodes for testing. On the other five heterophilic graphs, we leverage the sparse splitting (Chien et al., 2021) with 2.5%/2.5%/95% samples respectively for training/validation/testing. The results are reported in Table 1, where the best results are bold, the underlined letters denote the second highest accuracy, and the “Improv.” column indicates the relative enhancement of our models compared to BernNet. We first observe that both SAF and SAF-$\epsilon$ substantially boosts its base model, BernNet, with gains reaching a notable 15.37%. This impressive enhancement is credited to their capacity to effectively exploit the task-beneficial information, which is implicitly encoded by spectral filtering in the spatial domain. This ability is particularly advantageous in contexts with limited supervision, where it allows effective leveraging of extra prior knowledge during training. Generally, our models outperforms competitors on all datasets except for Chameleon and Squirrel, where it still make considerable improvements on BernNet by 14.56% and 9.68% and maintains a second-place rank. Moreover, we can observe that SAF-$\epsilon$ generally delivers better results than SAF. This advantage arises from its thresholding sparsity, reducing noise in non-local graphs for more efficient graph learning. However, this enhancement also incurs higher computational costs, as illustrated in both Section 5 and Appendix F.5.
Table 2: Full-supervised node classification accuracy (%) $\pm$ 95% confidence interval.
@@
Improv. 6.77% 12.61% 2.13% 1.15% 1.27% 1.36% 1.37% 3.01%

### 08 Feb 2024 v3

@@ 4. Rethinking Spectral GNNs from the Spatial Perspective
4.1. Interplay of Spectral and Spatial Domains through the Lens of Graph Optimization
The graph signal denoising problem (Shuman et al., 2013) was initially leveraged in (Ma et al., 2021; Zhu et al., 2021) as a means to interpret GNNs with smoothness assumption, which yet does not always hold in certain real-world graph scenarios such as heterophily (Zhu et al., 2020b). Without loss of generality, in this work, we consider a more generalized graph optimization problem [^1] This problem was first introduced in (He et al., 2021) for theoretically grounded graph filters. However, in this study, we repurpose it as a bridge between spectral filtering and spatial aggregation.
(3) $\mathop{\arg\min}_{\mathbf{Z}}\ \mathcal{L}=\alpha\|\mathbf{X}-\mathbf{Z}\|^{2}_{F}+(1-\alpha)\cdot\mathrm{tr}(\mathbf{Z}^{T}\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z})$
@@ 4.1. Interplay of Spectral and Spatial Domains through the Lens of Graph Optimization
(4) $\mathbf{Z}^{*}=(\mathbf{I}+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\hat{\mathbf{L}}))^{-1}\mathbf{X}=g_{\psi}(\hat{\mathbf{L}})\mathbf{X}=\mathbf{U}g_{\psi}(\mathbf{\Lambda})\mathbf{U}^{T}\mathbf{X}.$
@@
(5) $\mathbf{Z}^{(k)}=\alpha\mathbf{X}+(1-\alpha)\hat{\mathbf{A}}^{\text{new}}\mathbf{Z}^{(k-1)},\quad k=1,2,\cdots,K$
Theoretical Interaction — the Adapted New Graph. With the non-negative constraint, it is evident that both spectral filtering and spatial aggregation effectively address the generalized graph optimization problem in Eq. (3), despite their distinctive forms and operation domains. Upon closer examination, we discover a compelling relationship between the graph filter $g_{\psi}(\lambda)$ in Eq. (4) and the new graph $\hat{\mathbf{A}}^{\text{new}}$ in Eq. (5), given $g_{\psi}(\lambda)=(1+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\mathbf{\lambda}))^{-1}$,
(6) $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\gamma_{\theta}(\hat{\mathbf{L}})=\mathbf{I}-\frac{\alpha}{1-\alpha}(g_{\psi}(\hat{\mathbf{L}})^{-1}-\mathbf{I})$
@@
What are the differences between $\hat{\mathbf{A}}^{\text{new}}$ and $g_{\psi}(\hat{\mathbf{L}})$? Whereas the former as the uncovered new graph elucidates the inherent spatial node relationships, the latter is a graph operation, primarily processing graph features within spectral domain. It is crucial to understand that $g_{\psi}(\hat{\mathbf{L}})$ may not result in a dense matrix, especially with fixed-order polynomial approximation. This is because it captures up to only a $K$-hop neighborhood, i.e., $g_{\psi}(\hat{\mathbf{L}})=\sum_{k=0}^{K}\psi_{k}P_{k}(\hat{\mathbf{L}})=\sum_{k=0}^{K}\omega_{k}\hat{\mathbf{A}}^{k}$, practically limiting spectral GNNs’ effective propagation range. In contrast, our newfound graph $\hat{\mathbf{A}}^{\text{new}}$ intrinsically enjoys a non-local property, as confirmed in the following section. Building upon this discovery, we further devise a framework to break domain barriers, overcoming the limitations of current spectral GNNs due to truncated polynomials (see details in Section 5).
4.2. In-depth Analysis of the adapted new graph
@@ 4.2.1. Non-locality
With the established non-negative constraint on graph filters, specifically $0<g_{\psi}(\lambda)\leq 1$, it becomes evident that the eigenvalues of $\mathbf{I}-g_{\psi}(\hat{\mathbf{L}})$ falls into the interval permitting Neumann series expansion, as shown in lemma 1 (Horn and Johnson, 2012). Building on this observation, we present a non-trivial property of the new graph in the following proposition (see proof in Appendix A.1).
Proposition 0 ().
Given adjacency matrix $\hat{\mathbf{A}}^{\text{new}}$ formulated in Eq. (6), the adapted new graph exhibits non-locality. Specifically, $\hat{\mathbf{A}}^{\text{new}}$ is expressible as an infinite series expansion of the original graph’s adjacency matrix $\hat{\mathbf{A}}$. Formally, we have $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\frac{\alpha}{1-\alpha}\sum_{t=1}^{\infty}(\mathbf{I}-\sum^{K}_{k=0}\pi_{k}\hat{\mathbf{A}}^{k})^{t}=\sum_{t=0}^{\infty}\phi_{t}\hat{\mathbf{A}}^{t}$ where $\pi_{k}$ and $\phi_{t}$ refer to the constant coefficients computed from $\{\psi_{0},\psi_{1},...,\psi_{K}\}$ in distinct ways.
This proposition implies that the new graph engenders immediate links between nodes that originally necessitate multiple hops for connection. To further underpin this theoretical claim, we analyze the general connection status on the new graph by BernNet (He et al., 2021), a spectral GNN adhering to the non-negative constraint. From Figure 1, it is apparent that nodes originally separated by multiple hops achieve direct connections in the new graph.
4.2.2. Signed Edge Weights — Discerning Label Consistency
Upon further scrutinizing the adapted new graph, we make a notable discovery that it readily accommodates both positive and negative edge weights. A more granular analysis in Figure 2 reveals that a considerable portion of positive edge weights are assigned to the same-class node pairs, enhancing graph homophily (exemplified by edge homophily ratio (Zhu et al., 2020b)). Conversely, edges parameterized with negative weights tend to bridge nodes with different labels. These findings demonstrate the newfound graph’s adeptness in discerning label consistency among nodes. To theoretically explain this phenomenon, we further present the proposition below (see proof in Appendix A.2):
Proposition 0 ().
Let $\mathbf{Z}^{*}$ be the node representations optimized by Eq. (3). For $\mathbf{Z}^{*}$ to be effective in label prediction, it is a necessary condition that $\hat{\mathbf{A}}^{\text{new}}$ accommodates both positive and negative edge weights s.t. for any node pairs $v_{i},v_{j}\in\mathcal{V}$, $\hat{A}^{\text{new}}_{i,j}>0$ if $\mathbf{y}_{i}=\mathbf{y}_{j}$ and $\hat{A}^{\text{new}}_{i,j}<0$ if $\mathbf{y}_{i}\neq\mathbf{y}_{j}$.
@@ 4.2.2. Signed Edge Weights — Discerning Label Consistency
5. Spatially Adaptive Filtering Framework
Building on our discoveries, we re-evaluate the state-of-the-art spectral GNNs and put forth a paradigm-shifting framework, Spatially Adaptive Filtering (SAF), for joint exploitation of graph-structured data across both spectral and spatial domains (refer to Figure 3). SAF leverages the adapted new graph by spectral filtering for an auxiliary non-local aggregation, addressing enduring challenges in GNNs related to long-range dependencies and graph heterophily.
5.1. Non-negative Spectral Filtering
The proposed SAF requires explicit computation of the newfound graph, as outlined in Eq. (6). This further necessitates the graph filter $g_{\psi}:\left[0,2\right]\to\mathbb{R}$ to satisfy the non-negative constraint from Eq. (3): $0\leq g_{\psi}(\lambda)\leq 1$. However, not all extant graph filters fulfill this prerequisite. For instance, the filter use by GCN (Kipf and Welling, 2017), $g_{\psi}(\lambda)=1-\lambda$, takes negative values when $\lambda>1$. In this research, we approximate the graph filter using Bernstein polynomials (Farouki, 2012), which are known for their non-negative traits (Powers and Reznick, 2000) and are essential in a preeminent spectral GNN, BernNet (He et al., 2021). For $g_{\psi}(\lambda)\leq 1$ part, we rescale Bernstein polynomials with the following proposition.
Proposition 0 ().
@@ 5.1. Non-negative Spectral Filtering
where $f_{\varphi}(\cdot)$, a two-layer MLP, maps $\mathbf{X}$ from $F$ to $C$ dimensions using 64 hidden units, and $\{\psi_{k}\}_{k=1}^{K}$ are non-negative learnable parameters. Note that SAF also permits alternative implementations such as using Chebyshev polynomials (Hammond et al., 2011; Defferrard et al., 2016) for graph filter learning, enhancing models like ChebNetII (He et al., 2022) (see details in Appendix E.4).
5.2. Non-local Spatial Aggregation
@@
Once acquiring a suitable spectral filter $\hat{g}_{\psi}(\lambda)$, we compute the adapted new graph as $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\tau(\mathbf{U_{m}}g_{\psi}(\mathbf{\Lambda_{m}})^{-1}\mathbf{U_{m}}^{T}-\mathbf{I})$ by Eq. (6) where $\tau=\frac{\alpha}{1-\alpha}$ is a scaling parameter and a partial eigendecomposition can be employed to obtain only $m$ extremal eigenvalues (Lanczos, 1950), producing a low-rank, robust structure for $\hat{\mathbf{A}}^{\text{new}}$. Equipped with this newfound graph, we proceed to perform non-local aggregation:
$\mathbf{Z}^{(l)}=(1-\eta)\mathbf{Z}^{(0)}+\eta\hat{\mathbf{A}}^{\text{new}}\mathbf{Z}^{(l-1)},\quad l=1,2,\cdots,L$
@@
where $\eta$ refers to the update rate and $\mathbf{Z}^{(0)}=f_{\varphi}(\mathbf{X})$. The iteratively aggregated results are denoted as $\mathbf{Z}_{a}$. Recognizing the potential noise from the non-local nature of $\hat{\mathbf{A}}^{\text{new}}$, we apply a sparsification technique, leveraging a positive threshold $\epsilon$, and retain only essential elements outside the $\left[-\epsilon,\epsilon\right]$ interval. For clarity, this refined model is referred to as SAF-$\epsilon$.
Figure 3. Illustration of the proposed SAF framework, where varying node colors represent different node labels.
5.3. Node-wise Prediction Amalgamation
@@
To leverage information from different graph domains, we employ an attention mechanism, allowing nodes to determine the importance of each space. This mechanism produces pairwise weights for a nuanced amalgamation during prediction. Specifically, the weight pair is computed as $\boldsymbol{\kappa}_{f}=\text{Sigmoid}(\mathcal{P}_{f}(\mathbf{Z}_{f}))$, $\boldsymbol{\kappa}_{a}=\text{Sigmoid}(\mathcal{P}_{a}(\mathbf{Z}_{a}))$ where $\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\in\mathbb{R}^{N}$ contain the weights for each node, and $\mathcal{P}_{f}(\cdot)$ and $\mathcal{P}_{a}(\cdot)$ are two different mappings from $\mathbb{R}^{C}$ to $\mathbb{R}$. For simplicity, we implement them using two one-layer MLPs. Given domain predictions $\mathbf{Y}_{f},\mathbf{Y}_{a}\in\mathbb{R}^{C}$, the final model prediction is attained as
(7) $\mathbf{Y}=\mathrm{diag}(\boldsymbol{\kappa}_{f})\cdot\mathbf{Y}_{f}+\mathrm{diag}(\boldsymbol{\kappa}_{a})\cdot\mathbf{Y}_{a}$
@@
where a normalization $\left[\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\right]\leftarrow\frac{\left[\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\right]}{\max\{\|\{\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\}\|_{1},\delta\}}$ is performed beforehand to maintain $\boldsymbol{\kappa}_{f}+\boldsymbol{\kappa}_{a}=\mathbf{1}$ with small value $\delta$ preventing zero division. Similar schemes can be founded in works (Sagi and Rokach, 2018; Wang et al., 2020; Zhu et al., 2020a).
@@
Table 1. Semi-supervised node classification accuracy (%) $\pm$ 95% confidence interval.
@@
Improv.333 Improv. indicates the relative improvement of SAF over its base model, BernNet (He et al., 2021). For alternative implementation using ChebNetII (He et al., 2022), please refer to Appendix E.4. 14.56% 9.68% 15.37% 13.99% 3.97% 1.62% 1.86% 0.68%
Table 2. Full-supervised node classification accuracy (%) $\pm$ 95% confidence interval.
@@
Improv.333 Improv. indicates the relative improvement of SAF over its base model, BernNet (He et al., 2021). For alternative implementation using ChebNetII (He et al., 2022), please refer to Appendix E.4. 6.77% 12.61% 2.13% 1.15% 1.27% 1.36% 1.37% 3.01%
5.4. Complexity Analysis
SAF augments spectral GNNs with non-local aggregation and node-wise amalgamation. The first part entails creating a new graph and information propagation. In SAF-$\epsilon$, these two steps are separated by sparsification, culminating in $\mathcal{O}(N^{3}+N^{2}+nnz(\hat{\mathbf{A}}^{\text{new}})d)$ complexity, where $nnz$ denotes non-zero element count. Conversely, SAF, viewing non-local aggregation holistically, can reduce complexity to $\mathcal{O}(2dN^{2}+dN)$ when $d\ll N$. For node-wise amalgamation, its parallelizable nature ensures computational efficiency. Our method also requires a precomputation of eigendecomposition [^2] For more on its modern applications and discussions, please refer to Appendix B (including Specformer (Bo et al., 2023) & FE-GNN (Sun et al., 2023) featured in our experiments) and E.5. which, naively complex at $\mathcal{O}(N^{3})$, can be reduced to $\mathcal{O}(m^{2}+nnz(\hat{\mathbf{L}})m)$ using Lanczos method (Lanczos, 1950) for larger graphs with $m\ll N$ denoting iterative steps and is reusable for both training and inference. We present empirical studies on both time and space overheads in Appendix E.5.
6. Experiments
6.1. Datasets and Experimental Setup
6.1.1. Datasets
We evaluate models over 13 real-world datasets from various domains. These include three well-known homophilic graphs: Cora, Citeseer, and Pubmed (Sen et al., 2008), five commonly used heterophilic graphs: Chameleon, Squirrel (Rozemberczki et al., 2021), Cornell, Texas (Pei et al., 2020), and Actor (Tang et al., 2009), as well as five recently introduced benchmarks: Minesweeper (synthetic graph), Tolokers (crowdsourcing platform worker network), Amazon-ratings (product co-purchasing), Roman-empire (word dependency graph) (Platonov et al., 2023b), and Penn94 (Lim et al., 2021) (social network). Detailed statistics are provided in Appendix C.
6.1.2. Baselines
We compare SAF with 22 models: (1) MLP; (2) Basic GNNs: GCN (Kipf and Welling, 2017) & APPNP (Gasteiger et al., 2019); (3) Spectral GNNs: ARMA (Bianchi et al., 2021), GPR-GNN (Chien et al., 2021), BernNet (He et al., 2021), ChebNetII (He et al., 2022), JacobiConv (Wang and Zhang, 2022), Specformer (Bo et al., 2023), LON-GNN (Tao et al., 2023) & OptBasisGNN (Guo and Wei, 2023b); (4) Spatial GNNs: GCNII (Chen et al., 2020), PDE-GCN (Eliasof et al., 2021), GEN (Wang et al., 2021), NodeFormer (Wu et al., 2022), GloGNN++ (Li et al., 2022) & MGNN (Cui and Wei, 2023); (5) Unified GNNs: GNN-LF (Zhu et al., 2021), GNN-HF (Zhu et al., 2021), ADA-UGNN (Ma et al., 2021), FE-GNN (Sun et al., 2023) & ClenshawGCN (Guo and Wei, 2023a).
6.1.3. Setup
To follow (He et al., 2021, 2022; Wang and Zhang, 2022), we fix $K=10$. For each dataset, we perform a grid search to tune the hyper-parameters of all models. With the best hyper-parameters, we train models with Adam optimizer (Kingma and Ba, 2014) in 1,000 epochs using early-stopping strategy and a patience of 200 epochs, and report the mean classification accuracies with a 95% confidence interval on 10 random data splits. More experimental details can be found in Appendix D, and the codes will be made available if the paper could be accepted.
6.2. Overall Evaluation
6.2.1. Semi-supervised Node Classification.
In this task, we follow the experimental protocol established by (He et al., 2022) and compare our models with MLP, two basic GNNs, eight popular polynomial spectral GNNs, and five unified GNNs. For data splitting on homophilic graphs (Cora, Citeseer, and Pubmed), we apply the standard division (Yang et al., 2016) with 20 nodes per class for training, 500 nodes for validation, and 1,000 nodes for testing. On the other five heterophilic graphs, we leverage the sparse splitting (Chien et al., 2021) with 2.5%/2.5%/95% samples respectively for training/validation/testing. The results are reported in Table 1, where the best results are bold and the underlined letters denote the second highest accuracy. We first observe that both SAF and SAF-$\epsilon$ substantially boosts its base model, BernNet, with gains reaching a notable 15.37%. This impressive enhancement is credited to their capacity to effectively exploit the task-beneficial information, which is implicitly encoded by spectral filtering in the spatial domain. This ability is particularly advantageous in contexts with limited supervision, where it allows effective leveraging of extra prior knowledge during training. Generally, our models outperform competitors on all datasets except for Chameleon and Squirrel, where SAF maintains a second-place rank with considerable improvements on BernNet by 14.56% and 9.68%. In these cases, ChebNetII initially surpasses our model, yet, with more training samples, our SAF manages to beats it by margins of 3.93% and 6.28% (see Table 2). Moreover, it can be seen that SAF-$\epsilon$ averagely delivers better results than SAF, benefiting from its thresholding sparsity that reduces non-local noise for efficient graph learning. However, this enhancement also incurs higher computational costs, as illustrated in both Section 5 and Appendix E.5.
6.2.2. Full-supervised Node Classification
444 We borrow this terminology from (Chen et al., 2020; He et al., 2022) to denote 60%/20%/20% data splitting.